OpenStudy (anonymous):

Which method would be the simplest way to solve the system? y=1/2x 2x+3y=28 A. Graphing B. Substitution C. Elimination D. Distributive

OpenStudy (ranga):

Substitution. Replace y with x/2 in the second equation and you can quickly solve for x.

OpenStudy (ranga):

I don't think they are expecting you to solve the system but just which method is the simplest. But if they want a solution then: Put y = x/2 in second equation: 2x + 3x/2 = 28 multiply throughout by 2 4x + 3x = 56 7x = 56 x = 8 y = x/2 = 8/2 = 4 x = 8, y = 4

OpenStudy (anonymous):

Thank you! Would this one be substitution too: 7x+5y=19 -7x-2y=-16

OpenStudy (ranga):

Elimination would be the simplest for the second problem because if you add the two equations, the +7x and -7x will cancel out leaving you with just y which can be easily solved and then you can solve for x.

OpenStudy (anonymous):

Okay! Thanks!